### v1

# Extending Chevalley’s Theorem: A Topological Characterization of Constructibility and Its Generalization Beyond Noetherian Spaces

Abstract.
We introduce the notion of a good map between topological spaces: a continuous map $f:X\to Y$ is good if for every non-empty irreducible locally closed subset $U\subseteq X$, there exists a non-empty open subset $W\subseteq Y$ such that $W\cap f(U)=W\cap\overline{f(U)}\neq\varnothing$.
In Noetherian spaces, this condition is equivalent to preserving constructible subsets (Theorem 2.3), giving a topological characterization of Chevalley’s theorem. Without the Noetherian assumption, the good property continues to make sense and serves as a reasonable generalization.
We establish basic properties of good maps and introduce a weaker variant, weak good maps. In algebraic geometry, affine space projections $\mathbb{A}_{S}^{n}\to S$ are good for any base scheme $S$ (Theorem 4.1). From this we obtain:
  1. (1)
    A generalization of Chevalley’s theorem for morphisms of finite type with Noetherian underlying topological spaces (Theorem 4.2);
  2. (2)
    The fact that every morphism locally of finite type is weak good (Proposition 4.4), yielding an elementary proof of Jacobson ascent (Corollary 4.5).
The good property is stable under formal completion, suggesting extensions to other non-Hausdorff geometries.
Key words and phrases:
Chevalley’s theorem, constructible sets, good maps, weak good maps, Jacobson schemes, formal completion
2020 Mathematics Subject Classification:
14A15, 54C10, 13E05
1. Introduction
Chevalley’s theorem states that a morphism of finite type between Noetherian schemes preserves constructible subsets. This paper explores the topological nature of this preservation property.
We define a topological condition called goodness for continuous maps (Definition 2.1). In Noetherian spaces, a map is good if and only if it preserves constructible subsets (Theorem 2.3). This provides a topological characterization of Chevalley’s theorem in the Noetherian setting.
Without the Noetherian hypothesis, the good property still makes sense. It can be viewed as a generalization of constructibility preservation to arbitrary topological spaces. We develop basic properties of good maps and introduce a more flexible variant, weak good maps.
The theory finds concrete application in algebraic geometry. A key fact is that affine space projections $\mathbb{A}_{S}^{n}\to S$ are good for any base scheme $S$ (Theorem 4.1). From this we derive two main results: a generalization of Chevalley’s theorem under different hypotheses (Theorem 4.2), and the fact that locally finite type morphisms are weak good (Proposition 4.4), leading to an elementary proof of Jacobson ascent.
Finally, we show the good property is stable under formal completion, providing a bridge to formal geometry and suggesting similar developments in other non-Hausdorff geometries.
2. Good Maps
2.1. Definition
Definition 2.1 (Good map).
A continuous map $f:X\to Y$ is good if for every non-empty irreducible locally closed subset $U\subseteq X$, there exists a non-empty open subset $W\subseteq Y$ such that
$W\cap f(U)=W\cap\overline{f(U)}\neq\varnothing,$
where $\overline{f(U)}$ is the closure of $f(U)$ in $Y$.
Remark 2.2.
The condition means $f(U)\cap W$ is closed in $W$ and non-empty. In Hausdorff spaces, irreducible subsets are singletons, so all continuous maps into Hausdorff spaces are trivially good. The concept is meaningful only for non-Hausdorff topologies.
2.2. Characterization in Noetherian Spaces
Recall that in a Noetherian topological space, a subset is constructible if it is a finite union of locally closed subsets.
We will use the following well‑known criterion for constructibility (see [1, Tag 053Z]).
Proposition 2.3 (Constructibility criterion).
Let $X$ be a Noetherian topological space. A subset $Y$ of $X$ is constructible if and only if for any irreducible closed subset $F$ of $X$ such that $Y\cap F$ is dense in $F$, the set $Y\cap F$ contains a nonempty open subset of $F$.
Theorem 2.4.
Let $f:X\to Y$ be a continuous map between Noetherian topological spaces. Then $f$ is good if and only if $f$ preserves constructible subsets.
Proof.
(Preserves constructible $\Rightarrow$ good) Let $U\subseteq X$ be a non‑empty irreducible locally closed subset. It is constructible, so $f(U)$ is constructible in $Y$. Let $F=\overline{f(U)}$, which is irreducible because $U$ is irreducible and $f$ continuous. Since $f(U)$ is constructible and dense in $F$, Proposition 2.3 gives a non‑empty open subset $W_{F}\subseteq F$ such that $W_{F}\subseteq f(U)$. Choose an open subset $W\subseteq Y$ with $W\cap F=W_{F}$. Then
$W\cap f(U)=W_{F}=W\cap\overline{f(U)}\neq\varnothing.$
Hence $f$ is good.
(Good $\Rightarrow$ preserves constructible) Let $C\subseteq X$ be constructible; we show $f(C)$ is constructible. By Proposition 2.3, it suffices to show: for any irreducible closed subset $F\subseteq Y$ such that $f(C)\cap F$ is dense in $F$, the intersection $f(C)\cap F$ contains a non‑empty open subset of $F$.
Let $D=C\cap f^{-1}(F)$, a constructible subset of $X$. Write $D=\bigcup_{i=1}^{n}U_{i}$ as a finite union of locally closed subsets. Then
$f(C)\cap F=f(D)=\bigcup_{i=1}^{n}f(U_{i}).$
Since the union is dense in the irreducible closed set $F$, some $f(U_{i})$ is dense in $F$. Write $U_{i}$ as a finite union of its irreducible components $U_{ij}$. Again by irreducibility of $F$, some $f(U_{ij})$ is dense in $F$. Now $U_{ij}$ is a non‑empty irreducible locally closed subset of $X$. Since $f$ is good, there exists a non‑empty open subset $W\subseteq Y$ such that
$W\cap f(U_{ij})=W\cap\overline{f(U_{ij})}\neq\varnothing.$
But $\overline{f(U_{ij})}=F$ by density, so $W\cap F\subseteq f(U_{ij})\subseteq f(C)\cap F$. Hence $f(C)\cap F$ contains the non‑empty open subset $W\cap F$ of $F$, as required. ∎
Thus, in Noetherian spaces, goodness characterizes exactly those maps that preserve constructibility.
2.3. Basic Properties
Proposition 2.5.
Let $f:X\to Y$ be good.
  1. (1)
    For any locally closed $S\subseteq X$, $f|_{S}:S\to Y$ is good.
  2. (2)
    If $f(X)\subseteq Y^{\prime}\subseteq Y$, then $f:X\to Y^{\prime}$ (subspace topology) is good.
Proof.
Straightforward from the definition. ∎
Recall that a topological space is sober if every non-empty irreducible closed subset has a unique generic point. Schemes are sober topological spaces.
Proposition 2.6 (Glueing good maps).
Assume $X$ and $Y$ are sober. If $\{X_{i}\}$ is a finite open cover of $X$ and each restriction $f|_{X_{i}}:X_{i}\to Y$ is good, then $f:X\to Y$ is good.
Proof.
Let $U\subseteq X$ be a non-empty irreducible locally closed subset. Since $X$ is sober, the closure $\overline{U}$ has a unique generic point $x\in U$. Because $\{X_{i}\}$ is an open cover, $x$ belongs to some $X_{i}$. In fact, if $X_{i}\cap U\neq\varnothing$, then $X_{i}\cap U$ is an open subset of the irreducible space $U$, hence dense in $U$ and contains $x$. Thus $x\in X_{i}$ for every $i$ with $X_{i}\cap U\neq\varnothing$.
For each such $i$, the set $U_{i}:=X_{i}\cap U$ is a non-empty irreducible locally closed subset of $X_{i}$. Since $f|_{X_{i}}$ is good, there exists a non-empty open subset $W_{i}\subseteq Y$ such that
$W_{i}\cap f(U_{i})=W_{i}\cap\overline{f(U_{i})}^{Y}\neq\varnothing.$
Because $f(x)\in f(U_{i})$, we have $f(x)\in W_{i}$. Set $W=\bigcap_{i:U_{i}\neq\varnothing}W_{i}$; this is a finite intersection of non-empty open sets containing $f(x)$, hence a non-empty open subset of $Y$.
For each such $i$, we have
$W\cap f(U_{i})=W\cap\overline{f(U_{i})}.$
Taking unions over $i$ gives
$W\cap f(U)=\bigcup_{i}W\cap f(U_{i})=\bigcup_{i}W\cap\overline{f(U_{i})}=W\cap\bigcup_{i}\overline{f(U_{i})}.$
Since the union is finite, $\bigcup_{i}\overline{f(U_{i})}=\overline{\bigcup_{i}f(U_{i})}=\overline{f(U)}$. Therefore,
$W\cap f(U)=W\cap\overline{f(U)}\neq\varnothing,$
showing that $f$ is good. ∎
3. Weak Good Maps
3.1. Definition
Definition 3.1 (Weak good map).
A continuous map $f:X\to Y$ is weak good if for every non-empty locally closed subset $U\subseteq X$, there exists a non-empty locally closed subset $V\subseteq Y$ such that $V\subseteq f(U)$.
Every good map is weak good: for irreducible $U$, take $V=W\cap\overline{f(U)}$ from Definition 2.1.
3.2. Properties
Proposition 3.2.
Weak good maps are closed under composition and are local on both source and target.
Proof.
Obvious from the definition. ∎
4. Applications in Algebraic Geometry
4.1. Affine Space Projections are Good
Theorem 4.1.
For any scheme $S$, the projection $\pi:\mathbb{A}_{S}^{n}\to S$ is good.
Proof.
Let $U\subseteq\mathbb{A}^{n}_{S}$ be a non‑empty irreducible locally closed subset. We must produce a non‑empty open subset $W\subseteq S$ such that
$W\cap\pi(U)=W\cap\overline{\pi(U)}\neq\varnothing.$
First replace $S$ by $\overline{\pi(U)}$ with the reduced scheme structure; thus we may assume $S$ is integral and $\pi(U)$ is dense in $S$. Shrink $S$ to an affine open neighborhood of its generic point, so $S=\operatorname{Spec}A$ is affine and integral.
Let $Z=\overline{U}\subset\mathbb{A}^{n}_{S}$ with the reduced induced scheme structure; then $Z$ is integral. Since $U$ is open in $Z$, we may shrink $U$ further to a non‑empty distinguished open subset of $Z$, hence $U=\operatorname{Spec}B$ is affine and integral. The morphism $U\to S$ corresponds to an injection $A\hookrightarrow B$, and $B$ is a finitely generated $A$-algebra because it is a localization of a quotient of $A[t_{1},\dots,t_{n}]$.
A relative version of Noether normalization yields a non‑zero element $s\in A$ and a finite injective homomorphism
$A_{s}[T_{1},\dots,T_{r}]\hookrightarrow B_{s}.$
Set $W_{0}=D(s)\subset S$. The finite injective homomorphism above implies that the composition
$U_{s}:=\pi^{-1}(W_{0})\cap U\longrightarrow\operatorname{Spec}A_{s}[T_{1},\dots,T_{r}]\longrightarrow W_{0}$
is surjective; hence $\pi(U_{s})=W_{0}$.
Because we have taken $S=\overline{\pi(U)}$, we have $W_{0}\subseteq\pi(U)$ and $W_{0}=W_{0}\cap\overline{\pi(U)}$. Choose an open subset $W\subseteq S_{\text{original}}$ such that $W\cap S=W_{0}$ (this is possible since $S$ is closed in the original $S_{\text{original}}$ with the reduced structure). Then
$W\cap\pi(U)=W_{0}=W\cap\overline{\pi(U)}\neq\varnothing,$
which is exactly the good condition for $\pi$. ∎
This is a basic but important fact: affine space projections satisfy the good condition without any assumption on $S$.
4.2. Generalized Chevalley Theorem
Remark 4.2.
The classical Chevalley theorem states that a morphism of finite type between Noetherian schemes preserves constructible subsets. The Stacks Project [1, Tag 054K] generalizes it to morphisms that are quasi‑compact and locally of finite presentation, without any Noetherian assumption on the schemes themselves. Our Theorem 4.3 provides a different generalization: we keep the finite type hypothesis (rather than locally of finite presentation) but require only that the underlying topological spaces be Noetherian. This allows, for example, closed immersions defined by infinitely generated ideals, as long as the topological spaces remain Noetherian.
Theorem 4.3 (Generalized Chevalley Theorem).
Let $f:X\to Y$ be a morphism of schemes such that:
  1. (1)
    $X$ and $Y$ have Noetherian underlying topological spaces,
  2. (2)
    $f$ is of finite type.
Then $f$ is good. In particular, $f$ preserves constructible subsets.
Proof.
By Theorem 2.4, it suffices to show that $f$ preserves constructible subsets. Moreover, because any constructible set is a finite union of locally closed sets, it is enough to prove that for every non‑empty irreducible locally closed subset $U\subseteq X$, the image $f(U)$ is constructible in $Y$.
Since $Y$ is Noetherian, we can choose a finite affine open cover $\{Y_{i}\}$ of $Y$. Because $f$ is of finite type, each preimage $f^{-1}(Y_{i})$ can be covered by finitely many affine opens $X_{ij}$ such that $\mathcal{O}_{X}(X_{ij})$ is a finitely generated algebra over $\mathcal{O}_{Y}(Y_{i})$. Write $U_{ij}=U\cap X_{ij}$; each non‑empty $U_{ij}$ is an irreducible locally closed subset of $X_{ij}$. As $U=\bigcup_{i,j}U_{ij}$ is a finite union, it is enough to show that each $f(U_{ij})$ is constructible.
Now consider the restriction $f|_{X_{ij}}:X_{ij}\to Y_{i}$. Since $X_{ij}=\operatorname{Spec}B_{ij}$ and $Y_{i}=\operatorname{Spec}A_{i}$ with $B_{ij}$ finitely generated over $A_{i}$, we can choose a surjection $A_{i}[t_{1},\dots,t_{n}]\twoheadrightarrow B_{ij}$, which corresponds to a closed immersion $X_{ij}\hookrightarrow\mathbb{A}^{n}_{Y_{i}}$. Thus $f|_{X_{ij}}$ factors as
$X_{ij}\hookrightarrow\mathbb{A}^{n}_{Y_{i}}\xrightarrow{\pi}Y_{i},$
where the first map is a closed immersion (hence a closed map) and $\pi$ is the projection. By Theorem 4.1, $\pi$ is good; by Proposition 2.5(1), its restriction to the closed subscheme $X_{ij}$ remains good. In Noetherian spaces good maps preserve constructibility (Theorem 2.4). Therefore $f|_{X_{ij}}$ preserves constructibility. In particular, $f(U_{ij})$ is constructible in $Y_{i}$, hence also in $Y$ because constructibility is local.
Finally, $f(U)=\bigcup_{i,j}f(U_{ij})$ is a finite union of constructible sets, thus constructible. Hence $f$ preserves constructible subsets, and by Theorem 2.4 it is good. ∎
Remark 4.4.
This generalizes Chevalley’s theorem under different hypotheses than the Stacks Project version. For example, it applies to closed immersions defined by infinitely generated ideals, provided the topological spaces are Noetherian.
4.3. Locally Finite Type is Weak Good
Proposition 4.5.
Every morphism of schemes locally of finite type is weak good.
Proof.
The property of being weak good is local on source and target (Proposition 3.2), so we may assume $X=\operatorname{Spec}B$ and $Y=\operatorname{Spec}A$ are affine, with $B$ a finitely generated $A$-algebra. Choose a surjection $A[t_{1},\dots,t_{n}]\twoheadrightarrow B$, corresponding to a closed immersion $i:X\hookrightarrow\mathbb{A}^{n}_{Y}$. Let $\pi:\mathbb{A}^{n}_{Y}\to Y$ be the projection. Then $f=\pi\circ i$.
By Theorem 4.1, $\pi$ is good. Since $i$ is a closed immersion, $X$ is a closed (hence locally closed) subset of $\mathbb{A}^{n}_{Y}$. By Proposition 2.5(1), the restriction $\pi|_{X}:X\to Y$ is good. Every good map is weak good, therefore $f=\pi|_{X}$ is weak good. ∎
Corollary 4.6 (Jacobson ascent).
If $f:X\to Y$ is locally of finite type and $Y$ is Jacobson, then $X$ is Jacobson.
Proof.
Recall that a topological space is Jacobson iff every non‑empty locally closed subset contains a closed point.
Let $U\subseteq X$ be a non‑empty locally closed subset. By Proposition 4.5, $f$ is weak good, so there exists a non‑empty locally closed subset $V\subseteq Y$ with $V\subseteq f(U)$. Since $Y$ is Jacobson, $V$ contains a closed point $y$.
The fibre $f^{-1}(y)\cap U$ is non‑empty because $y\in f(U)$. Moreover, $f^{-1}(y)$ is locally closed in $X$ (it is the inverse image of the closed point $y$), hence $f^{-1}(y)\cap U$ is locally closed in $X$. As $f$ is locally of finite type, the fibre $f^{-1}(y)$ is a scheme of finite type over the residue field $\kappa(y)$. Schemes of finite type over a field are Jacobson; therefore the non‑empty locally closed subset $f^{-1}(y)\cap U$ contains a point $x$ that is closed in $f^{-1}(y)$.
Because $y$ is closed in $Y$ and $f$ is locally of finite type, the set $f^{-1}(y)$ is closed in $X$; moreover, a point that is closed in the closed subset $f^{-1}(y)$ is also closed in $X$. Thus $x$ is closed in $X$, and $U$ contains the closed point $x$, proving that $X$ is Jacobson. ∎
5. Outlook
5.1. Formal Completions
For schemes, the good property behaves well under formal completion along closed subschemes, providing a bridge to formal geometry. More precisely:
Proposition 5.1.
Let $f:X\to Y$ be a (weak) good morphism of schemes, $X^{\prime}\subseteq X$ and $Y^{\prime}\subseteq Y$ closed subschemes with $f(X^{\prime})\subseteq Y^{\prime}$. Then the induced map on formal completions $\widehat{f}:\widehat{X}_{X^{\prime}}\rightarrow\widehat{Y}_{Y^{\prime}}$ is (weak) good.
Proof.
Consider the restriction $f|_{X^{\prime}}:X^{\prime}\to Y^{\prime}$. Since $X^{\prime}$ is a closed (hence locally closed) subscheme of $X$, by Proposition 2.5(1) the map $f|_{X^{\prime}}$ is (weak) good. Moreover, because $f(X^{\prime})\subseteq Y^{\prime}$, we may view $f|_{X^{\prime}}$ as a map $X^{\prime}\to Y^{\prime}$ and apply Proposition 2.5(2) to see that it remains (weak) good when regarded as a map into the subspace $Y^{\prime}$.
The underlying topological space of the formal completion $\widehat{X}_{X^{\prime}}$ is homeomorphic to the underlying topological space of $X^{\prime}$; similarly $\widehat{Y}_{Y^{\prime}}$ is homeomorphic to $Y^{\prime}$. Under these homeomorphisms, the map $\widehat{f}$ corresponds precisely to $f|_{X^{\prime}}:X^{\prime}\to Y^{\prime}$. Because (weak) goodness is a purely topological condition, $\widehat{f}$ inherits the (weak) good property from $f|_{X^{\prime}}$. ∎
5.2. A General Pattern
The applications follow a pattern:
  1. (1)
    Find a natural class of maps that are good (e.g., affine space projections).
  2. (2)
    Use properties of good maps to show other maps are good or weak good.
  3. (3)
    Derive geometric consequences.
This suggests similar developments in other non-Hausdorff geometric categories, such as formal schemes or rigid analytic spaces, by finding appropriate analogues of affine space projections.